OpenStudy (anonymous):

Can someone explain : Derivative y = sqrt(ln(6x-1))

OpenStudy (anonymous):

chain rule for this one

OpenStudy (anonymous):

\[(f\circ g)'(x)=f'(g(x))\times g'(x)\] with \[f(x)=\sqrt{x}, f'(x)=\frac{1}{2\sqrt{x}},g(x)=\ln(6x-1), g'(x)=\frac{6}{6x-1}\]

OpenStudy (anonymous):

more simply put, the derivative of the square root of something is one over two square root of something, times the derivative of something

OpenStudy (anonymous):

Which is considered the outside, and the inside? \[1/2(1/u)\]? or is everything inside the square root considered the inside so \[1/2(u) \] Where u = ln(6x-1) or \[1/2(1/u)\] where u = 6x-1

OpenStudy (anonymous):

I responded without refreshing... alright so I would sub in g(x) \[\frac{ 1 }{ 2\sqrt{\ln(6x-1)} }\] Then multiple g'(x) by the f'(g(x) above?

OpenStudy (anonymous):

No updates, bad load times, can't respond... If this ever goes through this is what I figured \[\frac{ 1 }{ 2\sqrt{\ln(6x-1)} }*\frac{ 6 }{ 6x-1 }\] \[\frac{ 6 }{ 12x-2\sqrt{\ln(6x-1}}\] Final : \[\frac{ 3 }{ 12x-\sqrt{\ln(6x-1)} }\]